# Multi-Agent Reinforcement Learning

### v2

@@
(A Partial Survey of) Decentralized, Cooperative
Multi-Agent Reinforcement Learning
@@
CTE methods assume centralization during training and execution (e.g., with fast, free and perfect communication) and have the most information during execution. That is, the actions of each agent can depend on the information from all agents. As a result, a simple form of CTE can be achieved by using a single-agent RL method with the centralized action and observation spaces (maintaining a centralized action-observation history for the partially observable case). CTE methods can potentially outperform the decentralized execution methods (since they allow centralized control) but are less scalable as the (centralized) action and observation spaces can scale exponentially with the number of agents. CTE is typically only used in the cooperative MARL case since centralized control implies cooperation. CTDE methods are perhaps the most common as they can use centralized information during training but execute in a decentralized manner—using only information available to that agent during execution. As a result, they can be more scalable than CTE methods, do not require communication during execution, and can often perform well. CTDE fits most naturally with the cooperative case, but can be potentially applied in competitive or mixed settings depending on what information is assumed to be observed. Decentralized training and execution methods make the fewest assumptions and are often simple to implement. In fact, as I’ll discuss, any single-agent RL method can be used for DTE by just letting each agent learn separately. Of course, there are pros and cons to such approaches as we discuss below. It is worth noting that DTE is required if no offline coordination is available. That is, if all agents must learn during online interactions without prior coordination, learning and execution must both be decentralized. DTE methods can be applied in cooperative, competitive, or mixed cases but this text will focus on the cooperative MARL case.
@@
It is worth noting that this survey is partial in two respects. First, it is from my viewpoint. Second, it does not cover all work in decentralized, cooperative MARL. This subarea has less work than subareas such as CTDE but it is still more extensive than presented here. I have included work that I believe is important for understanding the main concepts in the subarea and apologize for those that I have omitted.
@@ 1 The cooperative MARL problem: The Dec-POMDP
Formally, a Dec-POMDP is defined by tuple $\langle\mathbb{I},\mathbb{S},\{{\mathbb{A}}_{i}\},T,R,\{{\mathbb{O}}_{i}\},O,\mathcal{H},\gamma\rangle$, where
  • $\mathbb{I}$ is a finite set of agents;
  • $\mathbb{S}$ is a set of states with designated initial state distribution ${{{b}}_{0}}$;
  • ${\mathbb{A}}_{i}$ is a set of actions for each agent $i$ with $\mathbf{\mathbb{A}}=\times_{i}{\mathbb{A}}_{i}$ the set of joint actions;
@@
  • ${\mathbb{O}}_{i}$ is a set of observations for each agent, $i$, with $\mathbf{\mathbb{O}}=\times_{i}{\mathbb{O}}_{i}$ the set of joint observations;
@@
where $P(s|\mathbf{h}{},{{{b}}_{0}})$ is the probability of state $s$ after observing joint history $\mathbf{h}{}$ starting from state distribution ${{{b}}_{0}}$ and $\mathbf{\pi}(\mathbf{h}{})$ is the joint action taken at the joint history. Also, $\mathbf{h}{}\mathbf{\pi}(\mathbf{h}{})\mathbf{o}$ represents $\mathbf{h}{}^{\prime}$, the joint history after taking joint action $\mathbf{\pi}(\mathbf{h}{})$ in joint history $\mathbf{h}{}$ and observing joint observation $\mathbf{o}$. In the RL context, algorithms do not iterate over states and observations to explicitly calculate this expectation but approximate it through sampling. For the finite-horizon case, $V^{\mathbf{\pi}}(\mathbf{h}{})=0$ when the length of $\mathbf{h}{}$ equals the horizon $\mathcal{H}$, showing the value function includes the time step from the history. The discount factor, $\gamma$, is typically set to 1 in the finite-horizon case and $\gamma\in[0,1)$ in the infinite-horizon case ($\mathcal{H}=\infty$) [^2] The episodic case (Sutton and Barto, 2018) is typically indefinite-horizon with termination to a set of terminal states with probability 1 and the infinite-horizon case is sometimes called ‘continuing.’.
@@
It is somewhat popular to define policies that depend only on a single observation rather than the whole observation history. That is, ${\pi}_{i}:{\mathbb{O}}_{i}\to{\mathbb{A}}_{i}$, rather than ${\mathbb{H}}_{i}\to{\mathbb{A}}_{i}$. This type of policy is ofter referred to as reactive or Markov since it just depends on (or reacts from) the past observation. These reactive policies are typically not desirable since they can be arbitrarily worse than policies that consider history information (Murphy, 2000) but can perform well or even be optimal in simpler subclasses of Dec-POMDPs (Goldman and Zilberstein, 2004). In general, reactive policies reduce the complexity of finding a solution (since many fewer policies need to be considered) but only perform well when the problem is not really partially observable.
@@ 2.1 Independent Q-learning (IQL)
While independent Q-learning ignores the other agents, it still depends on them in order to generate the joint reward $r$ and the local observation $o_{i}$ since both of these depend on the actions of all agents. That is, from agent $i$’s perspective, it is learning in a (single-agent) POMDP, as Algorithm 1 shows (the states of the POMDP would be the states of the Dec-POMDP plus the histories of the other agents).333 The idea of fixing other agent policies and then solving the resulting POMDP (either by planning methods (Nair et al., 2003) or RL methods (Banerjee et al., 2012)) has had success and results in a best response for agent $i$. Furthermore, iterating over agents, $i$, fixing other agent policies, $\neg i$, and calculating best responses will lead to a local optimum of the Dec-POMDP (Banerjee et al., 2012). Unfortunately, such methods require coordination to communicate which agent’s turn it is to learn so they are not appropriate for the DTE case.
@@ 2.2.1 Distributed Q-learning
In the more general stochastic case, distributed Q-learning is not able to distinguish between environmental stochasticity and agent stochasticity. That is, if getting a high-valued reward is unlikely (due to reward or transition stochasticity) distributed Q-learning will assume it can be gotten with probability one using max rather than taking the expectation (over other agent policies and stochastic transitions) as in Equation 1. As a result, distributed Q-learning is overly optimistic in the stochastic case and may perform poorly. Learning and incorporating the probabilities in Equation 1 and then calculating the expectation over the next step Q-value could fix this problem but would assume agents observe the actions of other agents which is not the case for decentralized training.
@@ 2.2.3 Lenient Q-learning
While it is possible to decay $\beta$ towards $\alpha$ as learning continues, to be more optimistic at the beginning of learning (when a lot of exploration is happening) and still be robust to stochasticity as learning converges, lenient Q-learning (Wei and Luke, 2016) allows agents to adjust this leniency towards exploration on a history-action pair basis. Specifically, lenient Q-learning maintains a temperature, $T({h}_{i},{a}_{i})$, per history-action pair that adjusts the probability that a negative update at a particular history-action pair will be ignored. These temperature values allow history-action pairs that have not been visited often to not be updated yet, while frequently visited history-action pairs are updated as in IQL.
@@ 2.3.2 Independent DRQN (IDRQN)
19:        Perform gradient descent on parameters $\theta$ with the loss: $\big{(}y-{Q}_{i}^{\theta}({h}_{i},{a}_{i})\big{)}^{2}$

### v3

# An Introduction to Decentralized Training and Execution in Cooperative Multi-Agent Reinforcement Learning

@@
An Introduction to
Decentralized Training and Execution in
Cooperative Multi-Agent Reinforcement Learning
@@
CTE methods assume centralization during training and execution (e.g., with fast, free and perfect communication) and have the most information during execution. That is, the actions of each agent can depend on the information from all agents. As a result, a simple form of CTE can be achieved by using a single-agent RL method with the centralized action and observation spaces (maintaining a centralized action-observation history for the partially observable case). CTE methods can potentially outperform the decentralized execution methods (since they allow centralized control) but are less scalable as the (centralized) action and observation spaces scale exponentially with the number of agents. CTE is typically only used in the cooperative MARL case since centralized control implies cooperation. CTDE methods are perhaps the most common as they can use centralized information during training but execute in a decentralized manner—using only information available to that agent during execution. As a result, they can be more scalable than CTE methods, do not require communication during execution, and can often perform well. CTDE fits most naturally with the cooperative case, but can be potentially applied in competitive or mixed settings depending on what information is assumed to be observed. Decentralized training and execution methods make the fewest assumptions and are often simple to implement. In fact, as I’ll discuss, any single-agent RL method can be used for DTE by just letting each agent learn separately. Of course, there are pros and cons to such approaches as we discuss below. It is worth noting that DTE is required if no offline coordination is available. That is, if all agents must learn during online interactions without prior coordination, learning and execution must both be decentralized. DTE methods can be applied in cooperative, competitive, or mixed cases but this text will focus on the cooperative MARL case.
@@
It is worth noting that this is an introduction to the field of decentralized, cooperative MARL. This subarea has less work than subareas such as CTDE but it is still more extensive than presented here. I have included work that I believe is important for understanding the main concepts in the subarea and apologize for those that I have omitted.
@@ 1 The cooperative MARL problem: The Dec-POMDP
Formally, a Dec-POMDP is defined by tuple $\langle\mathbb{I},\mathbb{S},\{{\mathbb{A}}_{i}\},T,R,\{{\mathbb{O}}_{i}\},O,\mathcal{H},\gamma\rangle$. For simplicity, I define the finite version but it easily extends to the continuous case:
  • $\mathbb{I}$ is a finite set of agents of size $|\mathbb{I}|=n$;
  • $\mathbb{S}$ is a finite set of states with designated initial state distribution ${{{b}}_{0}}$; [^1] Some papers refer to other information such as joint observations or histories as ‘state.’ For clarity, I will only use this term to refer to the true underlying state.
  • ${\mathbb{A}}_{i}$ is a finite set of actions for each agent $i$ with $\mathbf{\mathbb{A}}=\times_{i}{\mathbb{A}}_{i}$ the set of joint actions;
@@
  • ${\mathbb{O}}_{i}$ is a finite set of observations for each agent, $i$, with $\mathbf{\mathbb{O}}=\times_{i}{\mathbb{O}}_{i}$ the set of joint observations;
@@
I also include both the horizon and discount in the definition but, typically, only one is used. When $\mathcal{H}$ is finite, $\gamma$ can be set to 1 and when $\mathcal{H}=\infty$, $\gamma\in[0,1)$. [^2] The episodic case (Sutton and Barto, 2018) is typically indefinite-horizon with termination to a set of terminal states with probability 1 and the infinite-horizon case is sometimes called ‘continuing.’
@@
where $P(s|\mathbf{h}{},{{{b}}_{0}})$ is the probability of state $s$ after observing joint history $\mathbf{h}{}$ starting from state distribution ${{{b}}_{0}}$ and $\mathbf{\pi}(\mathbf{h}{})$ is the joint action taken at the joint history. Also, $\mathbf{h}{}\mathbf{\pi}(\mathbf{h}{})\mathbf{o}$ represents $\mathbf{h}{}^{\prime}$, the joint history after taking joint action $\mathbf{\pi}(\mathbf{h}{})$ in joint history $\mathbf{h}{}$ and observing joint observation $\mathbf{o}$. In the RL context, algorithms do not iterate over states and observations to explicitly calculate this expectation but approximate it through sampling. For the finite-horizon case, $V^{\mathbf{\pi}}(\mathbf{h}{})=0$ when the length of $\mathbf{h}{}$ equals the horizon $\mathcal{H}$, showing the value function includes the time step from the history.
@@
It is somewhat popular to define policies that depend only on a single observation rather than the whole observation history. That is, ${\pi}_{i}:{\mathbb{O}}_{i}\to{\mathbb{A}}_{i}$, rather than ${\mathbb{H}}_{i}\to{\mathbb{A}}_{i}$. This type of policy is ofter referred to as reactive or Markov since it just depends on (or reacts from) the past observation. These reactive policies are typically not desirable since they can be arbitrarily worse than policies that consider history information (Murphy, 2000) but can perform well or even be optimal in simpler subclasses of Dec-POMDPs (Goldman and Zilberstein, 2004). In general, reactive policies reduce the complexity of finding a solution (since many fewer policies need to be considered) but only perform well in limited cases such as when the problem is not really partially observable or when the observation history isn’t helpful for decision-making.
@@ 2.1 Independent Q-learning (IQL)
While independent Q-learning ignores the other agents, it still depends on them in order to generate the joint reward $r$ and the local observation $o_{i}$ since both of these depend on the actions of all agents. That is, from agent $i$’s perspective, it is learning in a (single-agent) POMDP, as Algorithm 1 shows (the states of the POMDP would be the states of the Dec-POMDP plus the histories of the other agents).444 The idea of fixing other agent policies and then solving the resulting POMDP (either by planning methods (Nair et al., 2003) or RL methods (Banerjee et al., 2012)) has had success and results in a best response for agent $i$. Furthermore, iterating over agents, $i$, fixing other agent policies, $\neg i$, and calculating best responses will lead to a local optimum of the Dec-POMDP (Banerjee et al., 2012). Unfortunately, such methods require coordination to communicate which agent’s turn it is to learn so they are not appropriate for the DTE case.
@@ 2.2.1 Distributed Q-learning
In the more general stochastic case, distributed Q-learning is not able to distinguish between environmental stochasticity and agent stochasticity. That is, if getting a high-valued reward is unlikely (due to reward or transition stochasticity) distributed Q-learning will assume it can be received deterministically using max rather than taking the expectation (over other agent policies and stochastic transitions) as in Equation 1. As a result, distributed Q-learning is overly optimistic in the stochastic case and may perform poorly. Learning and incorporating the probabilities in Equation 1 and then calculating the expectation over the next step Q-value could fix this problem but would assume agents observe the actions of other agents which is not the case for decentralized training.
@@ 2.2.3 Lenient Q-learning
While it is possible to adjust $\beta$ towards $\alpha$ as learning continues, to be more optimistic at the beginning of learning (when a lot of exploration is happening) and still be robust to stochasticity as learning converges, lenient Q-learning (Wei and Luke, 2016) allows agents to adjust this leniency towards exploration on a history-action pair basis. Specifically, lenient Q-learning maintains a temperature, $T({h}_{i},{a}_{i})$, per history-action pair that adjusts the probability that a negative update at a particular history-action pair will be ignored. These temperature values allow history-action pairs that have not been visited often to not be updated yet, while frequently visited history-action pairs are updated as in IQL.
@@ 2.3.2 Independent DRQN (IDRQN)
19:        Perform gradient descent on parameters $\theta$ with learning rate $\alpha$ and loss: $\big{(}y-{Q}_{i}^{\theta}({h}_{i},{a}_{i})\big{)}^{2}$